The sequence is given by the formula X(n+1) = 4Xn + √((15Xn^2)+(2)). It is known that X2017 + X2023 = 1464. Find X2020.
1
Expert's answer
2020-12-17T18:44:33-0500

It is known that "X_{n+1}=4X_n+\sqrt{15X_n^2+2}" and "X_{2017}+X_{2023}=1464" .

Note that "X_{n+1}= 4X_n+\sqrt{15X_n^2+2}>4X_n+\sqrt{9X_n^2}=4X_n+3|X_n|\geq 4X_n-3X_n=X_n"

It means that this sequence is strictly increasing and "X_n\cancel{=}X_k" for all "n\cancel{=} k" .

Using the formula of the "n" th term, we get:

"X_{n+1}-4X_n=\sqrt{15X_n^2+2}"

"X_{n+1}^2-8X_nX_{n+1}+16X_n^2=15X_n^2+2"

"X_{n+1}^2-8X_nX_{n+1}+X_n^2=2"


If we consider "n=k" and "n=k-1" , we get:

"\begin{cases}

X_{k}^2-8X_{k-1}X_{k}+X_{k-1}^2=2
\\
X_{k+1}^2-8X_kX_{k+1}+X_k^2=2

\\
\end{cases}
\quad \Rightarrow X_{k+1}^2-X^2_{k-1}-8X_{k}(X_{k+1}-X_{k-1})=0"

or "(X_{k+1}-X_{k-1})(X_{k+1}+X_{k-1}-8X_k)=0" .

Therefore, "X_{k+1}+X_{k-1}-8X_k=0" for all "k" . (1)

Let us consider "k=2020:"

"X_{2021}+X_{2019}=8X_{2020}" .

Using formula (1) for "k=2021" and "k=2019" , we get:

"\frac{X_{2020}+X_{2022}}{8}+\frac{X_{2018}+X_{2020}}{8}=8X_{2020}"

"X_{2022}+X_{2018}=62X_{2020}"

Using formula (1) for "k=2022" and "k=2018", we get:

"\frac{X_{2021}+X_{2023}}{8}+\frac{X_{2017}+X_{2019}}{8}=62X_{2020}"

"(X_{2021}+X_{2019})+(X_{2023}+X_{2017})=496X_{2020}"

"8X_{2020}+1464=496X_{2020}"

"488X_{2020}=1464"

"X_{2020}=3"


Answer: "X_{2020}=3."